# From News Sharers to Post Viewers: How Topic Diversity and Conspiracy Theories Shape Engagement With Misinformation During a Health Crisis?

### v2

@@ 4. Empirical Results
4.1. Topic Diversity (RQ1 & RQ2)
@@
Summary statistics: Fig. 3(a) shows that false news claims (mean of 0.086) exhibit slightly higher topic diversity compared to true news claims (mean of 0.082, , ). Notably, the difference in topic diversity between true and false news items is magnified at the post level, where the mean topic diversity in false news posts is 0.113, compared to 0.079 in true news posts (Fig. 3(b), , ). Additionally, from claims to posts (Fig. 3(c)), topic diversity in false news items (mean of 0.031) increases more than that in true news items (mean of 0.012, , ). For false news items containing conspiracy theories (CTs), the topic diversity at the claim level has no statistically significant difference compared to the overall false news items (Fig. 3(a), , ). However, Fig. 3(c) indicates a higher proportion of the segment with most significant increases in topic diversity for CT news items, compared to false news items ( , ), which may lead to the significant difference in topic diversity between CT news and general false news at the post level (Fig. 3(b), , ).

### v3

@@ 4. Empirical Results
4.1. Topic Diversity (RQ1.1 & RQ2.1)
4.1.1. Descriptive statistics.
@@
We start by examining topic diversity from news claims to news posts (Figs. 3(a)–3(c)). Fig. 3(a) shows that false news claims (mean of 0.086) exhibit 4.9% higher topic diversity compared to true news claims (mean of 0.082, , ). This difference is significantly magnified (43.0% higher) at the post level, where the average topic diversity in false news posts is 0.113, compared to 0.079 in true news posts (Fig. 3(b), , ). Additionally, we examine the paired changes in topic diversity from claims to their corresponding posts (Fig. 3(c)), finding that topic diversity in false news items (increase of 0.031) increases 158.3% more than that in true news items (increase of 0.012, , ). For false news items containing conspiracy theories (CTs), the topic diversity at the claim level has no statistically significant difference compared to the overall false news items (Fig. 3(a), , ). However, Fig. 3(c) indicates that the most significant increases in topic diversity from claims to posts are more likely to be in CT news items, compared to overall false news items ( , ). This may lead to the significant difference in topic diversity between CT news and general false news at the post level (Fig. 3(b), , ).
4.1.2. Regression results.